OpenStudy (anonymous):

5n-2(n-2)=-11 what is n? i got -5 but not sure

OpenStudy (accessdenied):

Have you plugged it back into your original equation? 5*(-5) - 2(-5 - 2) = -11 This would be a good way to check. If you get -11 = -11, then n=-5 is true.

OpenStudy (accessdenied):

I think it looks good, though. 5n - 2(n - 2) = -11 5n- 2*n - 2*(-2) = -11 5n - 2n + 4 = -11 3n + 4 = -11 Solving this equation does seem to get you n= -5.